OpenStudy (anonymous):

the equation of the line passing through the point (-2, 4) and perpendicular to y = -one halfx + 4

OpenStudy (anonymous):

when two lines are perpendicular their slopes are related as m1 = -1/m2 here you have y = -one halfx + 4 the slope of this line is -one half = -1/2 = m1 as m1 = -1/m2 thus -1/2 = -1/m2 now find m2

OpenStudy (anonymous):

after you get m2 use point slope form to get that equation point slope form is y-y1 = m2(x-x1) here (x1,y1)=(-2,4) and m2 = ......

OpenStudy (anonymous):

y=2x+8?

OpenStudy (anonymous):

yes exactly you did it man